OpenStudy (anonymous):

4y^2-81x^2 factor show steps

OpenStudy (anonymous):

(2y+9x)(2y-9x)

OpenStudy (anonymous):

4y^2-81x^2 take the square root of all variables and numbers involved (4 and 81,x^2 and y^2) so... (2y-9x)(2y+9x)

OpenStudy (anonymous):

is there anything you dont understand?

OpenStudy (anonymous):

equation is of the form \[(a^2-b^2)=(a+b)(a-b)\] here, a=2y, b= 9x